# Attenuation Bias

### v2

@@ Abstract
Many core concepts in political science are latent and therefore can only be measured with error. Measurement error in a predictor attenuates slope coefficient estimates in regression, biasing them toward zero. We show that widely used strategies for correcting attenuation bias—including instrumental variables and the method of composition—are themselves biased, sometimes even more than simple regression ignoring the measurement error altogether. We derive appropriate bias correction methods using split‑sample measurement strategies. Our approach is modular and can be easily deployed with additive score, factor, or machine learning models, requiring no joint estimation while yielding consistent slopes under standard assumptions. Simulations and applications—political knowledge, democracy indices, and text‑based sentiment—show stronger relationships after correction, sometimes by 50%. Open‑source software implements the procedure. Results underscore that latent predictors demand tailored error correction; otherwise, conventional practice can exacerbate bias.
@@ Introduction
Many key concepts in political science, such as political knowledge, ideology, and democracy, are fundamentally unobservable and can only be estimated based on sets of observed indicators. A large body of research relies on measures of these latent variables as predictors in regression models to test theories and make empirical claims. For example, studies have employed proxies for political knowledge to explain patterns of political participation [11], used estimates of individuals’ ideology to predict vote choice and other political behaviors [46, 26], and used democracy indices to assess the relationship between democratic institutions and economic growth [1] or the risk of conflict [21].
@@
Because these latent variables must be estimated based on observed indicators, these measurements necessarily will include error. Latent variables also have no inherent scale, so researchers must impose identifying restrictions to pin these measurements down. This, combined with measurement error, produces coefficient estimates that are biased in different ways compared to the standard classical measurement error case with observable variables. While the problems of measurement error and attenuation bias are widely recognized [17, 12, 7, 37, 16], the additional challenges posed by latent predictors are often underemphasized in applied work. Indeed, in an original survey of all articles published in the three top substantive political science journals (APSR, AJPS, JOP as well as Political Analysis) between 2000 and 2023, we find that while a growing number of articles mention latent variables or measurement error, a small and relatively constant number discusses the two in combination (see Appendix VII).
@@
In this paper, we examine the issues associated with using estimates of latent variables as predictors and present methods to mitigate the resulting biases under specific assumptions. We build on previous work, including [2] and [5], while deriving important new results and introducing new tools that account for key features of latent variables and their estimation. We show that two commonly used techniques for adjusting regression estimates with latent predictors can and often do exacerbate attenuation bias, sometimes dramatically so. Most notably, the method of composition [47], which is the most prominent method used in this area, actually makes the attenuation bias due to measurement error worse than it would be if the measurement error had simply been ignored altogether.
@@
Building on split-sample methods [3], we use multiple estimates of latent traits based on disjoint sets of indicators to correct standard OLS and 2SLS estimators, which are each biased in opposing directions in the latent predictor case. The method is modular—compatible with additive/averaging scores, factor models, and machine‑learning approaches—and avoids the challenges of full joint estimation of measurement and outcome models (e.g., [9, 42]).
@@
We illustrate these methods with several applications, showing that correcting estimates for measurement error can sometimes result in substantively large changes, in some cases up to 50% in magnitude. But in other cases, measurement error may be mild enough that these corrections are smaller, perhaps even negligible. We also present theory and simulation results that demonstrate that while the naive OLS estimator (ignoring measurement error in the latent predictor) is biased toward zero and previously used methods for addressing measurement error can result in even more severely biased estimates, our corrected estimator dramatically reduces bias and produces estimates very similar to those obtained using the more cumbersome full joint estimation approach. Finally, we provide open-source software for implementing these methods, helping applied researchers account for measurement error when using latent predictors in practice.
1 Setup: Latent Predictors, Identification, and Notation
Our use of the term latent is meant to indicate not only that we do not have a true measure of the variable, but also that the variable is fundamentally unobservable and lacks an objective scale. For example, the values of variables such as knowledge or ideology can only be measured indirectly by examining indicators such as the number of factual questions a respondent answers correctly (for knowledge) or the number of policy questions a respondent gives the conservative response (for ideology). Of course, we could define the variable of interest in these cases as the proportion of questions answered correctly on a particular test (or the proportion of responses given in the conservative direction), which would mean that we do, in fact, have the variable itself measured. But it is likely that we actually care about some true underlying level of knowledge or of liberal-conservative ideology, and that the responses to these indicators are useful because they give us (imperfect) information about each respondent’s true value on these latent scales.
To put it more formally, we use the term latent variable to refer to a variable that is not directly observable, can only be estimated based on a set of stochastic indicators, and that does not have an objectively defined scale, so therefore must be measured on a scale that is defined by one or more identifying restrictions. Using these sorts of variables as predictors in regression models presents additional complications above and beyond the well-known issues with measurement error in (non-latent) predictors, as we discuss below. [^1] We do not consider here the important issue of whether a measurement strategy for a latent trait is theoretically related to the concept of interest (e.g., [36]). Instead, we focus on issues that arise conditional on the measurement strategy chosen by researchers.
It should be noted that, while data-rescaling is usually not a source of bias (see preceding reference to the Maximum Likelihood invariance principle), in this case, the dynamics are different because the scaling factors are data- and measurement error-dependent. [^2] We provide a tutorial outlining this intuition. See: GitHub.com/cjerzak/lpme-software/blob/main/tutorial/IntroTutorial.R?raw=true
Below, we briefly outline the additional complications that arise when using a predictor that is measured with error, is latent, and therefore has no objective scale. The key issue is that some scale must be assumed for the latent trait (i.e., we must pin down the variable’s meaning), both for the true unobserved variable $X$ and for the observed but measured with error variable, $\hat{X}$.
A commonly used identification restriction is that the true latent variable $X$ has mean 0 and variance 1, which identifies the scale in the common one-dimensional setting, except for direction (sign flips). [^3] See [40] for a formal discussion of identification in item response models. The direction of the scale is typically specified as well, with a corresponding identifying restriction imposed, e.g., that higher values of a democracy scale mean that a country is more democratic. In our discussion here, we adopt this typical identification restriction, assuming that the true latent variable $X$ has a mean of 0 and a variance of 1 within the sample. This identification restriction has the advantage of making the latent variable’s scale essentially a $z$-score, indicating how many standard deviations above or below the mean any value is, which can aid interpretability. The results below generally apply similarly to other possible identification restrictions, since changing the identification restrictions will simply impose an affine transformation on the estimated latent traits.
We begin by outlining some key definitions and introducing notation. (Where unambiguous, we omit unit subscripts, $i$.)
  • $X$ is the true (unobserved) latent trait, which is assumed to have mean 0 and standard deviation 1 across observations in the sample.
  • $\tilde{X}$ is the un-transformed estimated latent trait, which is assumed to be equal to the true latent trait plus some mean-zero homoskedastic random error, with the error being independent of the true latent trait. Formally, $\tilde{X}=X+U$ where $U\perp\!\!\!\perp X$, $\mathbb{E}[U]=0$ and $\textrm{Var}(U)=\sigma_{U}^{2}$ so that $\textrm{Var}(\tilde{X})=\textrm{Var}(X)+\textrm{Var}(U)=1+\sigma_{U}^{2}$.
  • $\hat{X}$ is the transformed (i.e., identified) set of estimates of the latent trait that is rescaled to have mean zero and standard deviation one within the sample. In other words, $\hat{X}$ is equal to $\tilde{X}$ minus the sample mean of $\tilde{X}$, all divided by the sample standard deviation of $\tilde{X}$.
Our central parameter of interest is the slope coefficient in a linear regression model predicting some dependent variable $Y$ with the latent trait $X$, which we can write as
@@ 1 Setup: Latent Predictors, Identification, and Notation
However, as described above, the fact that $X$ is a latent variable means that not only is it measured with error, but the scale of our measurements of $X$ is different in subtle ways. Figure 1 presents a visualization of how measurement error operates in this case where the true latent trait $X$ is assumed to have mean 0 standard deviation 1, and the observed estimates of the latent trait $\hat{X}$ are also rescaled to have mean 0 and variance 1.
Figure 1: Illustration of how measurement error operates in the classical case (“With Noise”) and in the case with identification restrictions (“With Noise & Normalization”); normalization is needed in order to fix the scale of the latent trait estimation, but that normalization is data- and measurement-error-dependent.
In the standard (non-latent variable) measurement error setup with an observable predictor, we would observe $\tilde{X}$, which is equal to the true $X$ plus measurement error, and could use it as our predictor. Importantly, note that under these standard assumptions $\tilde{X}$ is on the same scale as, but has larger variance than, the true latent trait $X$ since $\tilde{X}$ is equal to $X$ plus an error term that is independent of $X$. For example, if $X$ was a person’s true weight, we might have $\tilde{X}$ that is the reading from a scale that is equal to the true weight plus some random measurement error. In this case, $\tilde{X}$ would still be on the same scale as $X$ and, importantly, the variance of $\tilde{X}$ would be greater than the variance of $X$. In the case of a latent predictor, not only do we not observe the true variable $X$, but we also do not observe $\tilde{X}$. Instead, we must use the identified and transformed $\hat{X}$ as our predictor, which is rescaled to have the same variance as the true $X$. We can think of $\hat{X}$ as formed by taking the untransformed estimates $\tilde{X}$ and subtracting off the sample mean and dividing by the sample standard deviation. Note that we do not typically know what the variance of $\tilde{X}$ is without knowing the variance of the measurement error $\sigma^{2}_{U}$, so we cannot know what the scale of $\tilde{X}$ would be.
2 Why Attenuation Bias Differs with Latent Predictors
We consider a bivariate (for now) linear regression model predicting some dependent variable $Y$ with the true latent predictor $X$ or, alternatively, with $\tilde{X}$ or with $\hat{X}$, under standard linear model assumptions and assuming that the error term in the linear regression is independent of the measurement error $U$. We use the following notation:
  • $\beta_{X}$, our central parameter of interest, is the true slope coefficient on the (true) latent variable $X$ in the bivariate linear regression predicting $Y$
  • $\hat{\beta}_{X}$ is the least squares estimator of $\beta_{X}$ if we were to use the true $X$ (measured without error) as our predictor
  • $\hat{\beta}_{\tilde{X}}$ is the least squares estimator of $\beta_{X}$ if we were to use the un-transformed estimated latent trait $\tilde{X}$ as our predictor
  • $\hat{\beta}_{\hat{X}}$ is the least squares estimator of $\beta_{X}$ if we were to use the transformed (i.e., identified) latent trait estimates $\hat{X}$ as our predictor
We call $\hat{\beta}_{\hat{X}}$ the naive estimator of the slope since it ignores measurement error in the latent predictor and instead treats point estimates $\hat{X}$ as if they were the true values of the predictor. As mentioned above, in the latent predictor case we generally do not observe either the true latent variable $X$ or the un-transformed estimates $\tilde{X}$, so our discussions of $\hat{\beta}_{X}$ and $\hat{\beta}_{\tilde{X}}$ will be for purposes of illustration and comparison to $\hat{\beta}_{\hat{X}}$ and to the corrected estimators we propose below.
If we were able to use the true latent variable $X$ as our predictor, then the standard linear regression assumptions would ensure that our estimator $\beta_{X}$ of the slope coefficient is unbiased. Of course, the fact that the predictor is latent means that we can never observe it directly, and hence we can never measure it without error.
If, by contrast, we were to use the measured with error predictor $\tilde{X}$, which is on the same scale as the true latent variable $X$, as our predictor, then the slope coefficient in a bivariate regression would have the standard measurement error attenuation factor [17, 10]:
$\mathbb{E}[\hat{\beta}_{\tilde{X}}]=\frac{\sigma^{2}_{X}}{\sigma^{2}_{X}+\sigma^{2}_{U}}\;\beta_{X}=\frac{1}{1+\sigma^{2}_{U}}\;\beta_{X}=\lambda\;\beta_{X}$ (2)
since the identifying assumptions for $X$ include its true values having variance 1. Here $\lambda\equiv\nicefrac{{1}}{{(1+\sigma^{2}_{U})}}$ is called the attenuation factor. For example, if on the true latent scale, the measurement error variance were equal to the variance of the latent trait itself (i.e., $\sigma_{U}^{2}=1$ then the attenuation factor would be $\frac{1}{1+1}=\frac{1}{2}$ and the estimated coefficient when using $\tilde{X}$ as the predictor would be half of the size of the true coefficient in expectation.
But when $X$ is a latent variable, we cannot observe $\tilde{X}$. Instead, we observe a noisy but identified estimate of $X$ that has been standardized to have mean 0 and variance 1. Thus, we can think of $\hat{X}$ as being equal to $\frac{\tilde{X}}{\sqrt{1+\sigma_{U}^{2}}}$ (ignoring mean adjustment, which affects only the intercept). It follows that the naive estimator using the identified latent trait estimates is
$\mathbb{E}[\hat{\beta}_{\hat{X}}]=\frac{1}{\sqrt{1+\sigma_{U}^{2}}}\beta_{X}$ (3)
meaning that the attenuation in this case is less severe than in the observable case (since $0<\frac{1}{\sqrt{1+\sigma_{U}^{2}}}>\frac{1}{1+\sigma_{U}^{2}}<1$ for $\sigma_{U}^{2}>0$).
Figure 2: Attenuation factor for estimated slope coefficient as a function of measurement error variance for observable predictor measured with error ($Y$ regressed on $\tilde{X}$) and latent predictor measured with error ($Y$ regressed on $\hat{X}$).
Figure 2 shows the attenuation factor for the bivariate regression slope coefficient estimate as a function of the measurement error variance $\sigma_{U}^{2}$ for the cases of observable and latent predictors. In both cases, it is assumed that the true predictor has variance 1 (equivalently, one can think of the variance of the measurement error as being the ratio of the variance of the measurement error to the variance of the true predictor). When there is no measurement error, the attenuation factor is 1 in both cases. As measurement error increases, both attenuation factors drop below 1, meaning that the expected value of the slope coefficient estimate is smaller in magnitude than the true parameter value. In the latent setting, the scale of the measured-with-error predictor is transformed by the identifying restriction to have variance 1—the same variance as the true variable measured without error. This rescaling multiplies the estimated slope coefficient, making it larger in magnitude and thereby mitigating, but not eliminating, attenuation.
It follows from the different attenuation factors for the latent predictor setup we consider here that applying standard corrections for attenuation based on estimates of measurement error will over-adjust the coefficient, resulting in estimates that are inflated, i.e., biased away from zero. In other words, if one corrects for an attenuation factor of $\frac{1}{1+\sigma^{2}_{U}}$ when the true attenuation factor is $\frac{1}{\sqrt{1+\sigma^{2}_{U}}}$ then this would be an over-correction resulting in amplification bias in the estimated slope coefficient.
@@ 3 Existing Approaches and Their Limitations
The most common approach to dealing with measurement error in latent predictors has been to simply ignore the problem and use point estimates of the latent trait as the predictor without any further adjustment. In cases where measurement error is minimal, using this naive estimator may be reasonable. But with significant measurement error, slope estimates will be biased toward zero, sometimes dramatically so, when using standard regression estimates.
A large literature in econometrics and other areas has dealt with so-called errors-in-variables modeling, with the most prominent of these techniques being instrumental variables (see [18] for an overview of these approaches). As mentioned above, the method of composition has been prominently used in recent decades in political science and related literature. Finally, it is sometimes possible to use full joint estimation in which the measurement model for the latent trait and the regression predicting $Y$ with $X$ are estimated together in a unified model. This section briefly discusses each of these approaches, including their benefits and drawbacks.
3.1 Instrumental Variables (IV): Over‑correction under Latent Scaling
One common method for correcting for measurement error in a predictor is instrumental variables (IV). Although its use today is often motivated by causal identification, IV can also, under certain assumptions, produce consistent estimates of slope coefficients when a predictor is measured with error. [^4] Although we focus on measurement error here, note that in a causal framework, IV can be used to address measurement error as well as simultaneity and omitted variable bias under suitable assumptions (see [37]). In the standard observable predictor setting, if $X$ is measured with error following the setup above and we seek to estimate the slope coefficient on $X$ in a linear regression predicting some other variable $Y$, we can correct for the resulting attenuation bias by leveraging an instrument $Z$ with the following properties: (i) $\textrm{Cov}(X,Z)\neq 0$; (ii) $\textrm{Cov}(Z,U)=0$; (iii) $\textrm{Cov}(Z,\varepsilon)=0$. In other words, a valid instrument in this setting must be correlated with the true predictor $X$ and be uncorrelated with both the measurement error in $X$ and the error term in the regression predicting $Y$ with $X$.
Standard IV adjustment for measurement error, however, will over-correct in the case of latent predictors. This is because the attenuation factor is different with latent predictors as seen in Equation 3 above. We show below that using standard IV can actually increase bias, resulting in coefficients that are too large in magnitude.

### v3

@@ Abstract
Many core concepts in political science are latent and therefore can only be measured with error. Measurement error in a predictor attenuates slope coefficient estimates in regression, biasing them toward zero. We show that widely used strategies for correcting attenuation bias—including instrumental variables and the method of composition—are themselves biased when applied to latent regressors, sometimes even more than simple regression ignoring the measurement error altogether. We derive a correlation-based correction using split‑sample measurement strategies. Rather than assuming a particular estimation strategy for the latent trait, our approach is modular and can be easily deployed with a wide variety of latent trait measurement strategies including additive score, factor, or machine learning models, requiring no joint estimation while yielding consistent slopes under standard assumptions. Simulations and applications show stronger relationships after our correction, sometimes by as much as 50%. Open‑source software implements the procedure. Results underscore that latent predictors demand tailored error correction; otherwise, conventional practice can exacerbate bias.
@@ Introduction
Many key concepts in political science, such as political knowledge, ideology, and democracy, are fundamentally unobservable and can only be estimated based on sets of observed indicators. A large body of research relies on measures of these latent variables as predictors in regression models to test theories and make empirical claims. For example, studies have employed proxies for political knowledge to explain patterns of political participation [undefk], used estimates of individuals’ ideology to predict vote choice and other political behaviors [undefan, undefv], and used democracy indices to assess the relationship between democratic institutions and economic growth [undef] or the risk of conflict [undeft].
@@
Because latent variables are unobserved and must be inferred from indicators, any empirical proxy for them (e.g., an index score or latent‑trait estimate) inevitably contains measurement error. While the problems of measurement error and attenuation bias are widely recognized [undefp, undefl, undeff, undefae, undefo], the additional challenges posed by latent predictors are often underemphasized in applied work. Indeed, we conducted an original survey of all articles published in the three top substantive political science journals (APSR, AJPS, JOP as well as Political Analysis) between 2000 and 2023. We find that while a growing number of articles mention latent variables or measurement error, a small and relatively constant number discusses the two in combination (see Appendix VII).
@@
In this paper, we examine the bias introduced by using estimated latent variables as predictors and present methods to mitigate it under specific assumptions. We build on previous work, including [undefa] and [undefd], while deriving important new results and introducing new tools that account for key features of latent variables and their estimation. We show that two commonly used techniques for adjusting regression estimates with latent predictors can and often do exacerbate bias, sometimes dramatically so. Most notably, the method of composition [undefao], which is the most prominent method used in this area, typically makes the attenuation bias due to measurement error worse than it would be if the measurement error had simply been ignored altogether.
@@
Building on split-sample methods [undefb], we use multiple estimates of latent traits based on disjoint sets of indicators to correct standard OLS and 2SLS estimators, which are each biased in opposing directions in the latent predictor case. The method is modular—compatible with a wide range of latent trait estimation approaches, including additive/averaging scores, factor models, and machine‑learning approaches—and avoids the challenges of full joint estimation of measurement and outcome models (e.g., [undefh, undefai]).
@@
With this modular correction in hand, we first illustrate its consequences in applications, showing that accounting for latent-predictor measurement error can shift estimated relationships by as much as 50 in magnitude. But in other cases, measurement error may be mild enough that these corrections are smaller, perhaps even negligible. We also present theory and simulation results. These show that with latent predictors the standard OLS estimator (ignoring measurement error in the latent predictor) is biased toward zero, and that previously used methods for addressing measurement error can result in even more severely biased estimates. By contrast, our corrected estimator dramatically reduces bias and produces estimates very similar to those obtained using the more cumbersome full joint estimation approach. Finally, we provide open-source software—an R package, lpmec ( GitHub.com/cjerzak/lpmec-software )—for implementing these methods, helping applied researchers account for measurement error when using latent predictors in practice.
1 Measurement Error and Bias in Linear Regression
1.1 Standard Classical Measurement Error
Before considering the latent predictor case in depth, we briefly review the case of classical measurement error in the standard (non-latent variable) case. We define the standard (non-latent) measurement-error case as where the regressor $X$ has an intrinsic scale but is observed with additive noise $\tilde{X}=X+U$. Here, and throughout, our parameter of interest is the slope $\beta_{X}$ in a linear regression of an outcome $Y$ on a predictor $X$:
@@ 1.1 Standard Classical Measurement Error
We use $\beta_{X}$ to denote the true regression coefficient that we seek to estimate and $X$ is the true predictor variable, measured without error.
Under the the standard classical measurement error setup the true predictor $X$ is not directly observed, but we instead observe
$\tilde{X}=X+U,$ (2)
where $\mathbb{E}[U]=0$, $U$ is independent of both $X$ and $\epsilon$, and $\textrm{V}(U)=\sigma_{U}^{2}$ so that $\textrm{V}(\tilde{X})=\textrm{V}(X)+\textrm{V}(U)$. When estimating a linear regression using noisy $\tilde{X}$, rather than $X$ itself, as our predictor, the least squares estimate of $\beta_{X}$ is biased according to the standard measurement error attenuation factor [undefp, undefj]:
$\operatornamewithlimits{plim}\hat{\beta}_{\tilde{X}}=\frac{\sigma^{2}_{X}}{\sigma^{2}_{X}+\sigma^{2}_{U}}\;\beta_{X}=\lambda\;\beta_{X},$ (3)
where $\lambda\equiv\nicefrac{{\sigma^{2}_{X}}}{{(\sigma^{2}_{X}+\sigma^{2}_{U})}}$ is called the attenuation factor.
1.2 Measurement Error in Latent Predictors
When the predictor $X$ is latent and must be estimated (and then scaled for identification), regression coefficients are attenuated by a different factor than in the standard (observable‑predictor measurement‑error) case. To explore this, we use a simple framework to generalize from the standard measurement error case involving noisy $\tilde{X}$ above.
We remain agnostic about the measurement strategy: $X$ may be reflective (indicators are modeled as responses to $X$) or formative (indicators are combined to define $X$); our results do not depend on the specific strategy used (see [undefi]). Instead, we assume that a researcher already has some approach to estimating the latent variable $X$ based on a set of observed indicators WW. [^1] We do not consider whether a measurement strategy for a latent trait is theoretically related to the concept of interest (e.g., [undefad]). For example, latent variables such as knowledge or ideology can be measured by examining indicators such as the number of factual questions a respondent answers correctly (for knowledge) or the number of policy questions a respondent gives the conservative response (for ideology).
Regardless of the measurement procedure, we model the unstandardized estimate as $\tilde{X}=X+U$, where $U$ captures the estimation error induced by using indicators to infer $X$ rather than observing $X$ directly. We later show that our approach is likely to perform well even when some of the above assumptions about $U$ may be violated.
Importantly, unlike in the standard measurement error case, latent traits variable estimates must be also identified. That is, they must be anchored since they do not have inherent objective scales. We assume that the true latent trait has mean 0 and variance 1 for units in the sample, which identifies the scale in the common one-dimensional setting, except for direction (sign flips). [^2] See [undefag] for a formal discussion of identification in item response models. The direction of the scale is typically specified as well, with a corresponding identifying restriction imposed, e.g., that higher values of a democracy scale mean that a country is more democratic. $\beta_{X}$, our ultimate quantity of interest, is only interpretable relative to the scale of the latent predictor. Multiplying the latent variable by two, for example, would halve its regression coefficient. Therefore, we maintain the simple and interpretable mean 0, standard deviation 1 identification restriction for our proposed method (and for all others examined) to make comparisons more straightforward.
Crucially, this identification restriction applies both to the true latent trait (e.g., imposing $\mathbb{E}[X]=0$ and $\textrm{V}(X)=1$, conceptually) and to whatever point estimates of the latent trait are used in subsequent analyses (e.g., empirically standardizing estimates to have mean zero and variance one). This distinction between the latent scale and the scale of the identified estimates is often ignored in applied work, but it is central to understanding how measurement error operates in the case of latent predictors and to the attenuation factor we derive below.
A natural question is whether this identification-driven rescaling is inevitable; in fact, one could instead report scores on a more “classical” scale—so that their variance matches $\textrm{V}(\tilde{X})=\sigma_{X}^{2}+\sigma_{U}^{2}$—and then apply the usual attenuation correction. Doing so, however, requires knowing (or estimating) $\sigma_{U}^{2}$; and once $\sigma_{U}^{2}$ is available, combining this rescaling with the usual attenuation correction yields, in fact, the same coefficient adjustment as our proposed approach (as we show below).
Formally, we assume that our point estimates of a latent variable predictor are
$\hat{X}=\frac{\tilde{X}-\text{mean}(\tilde{X})}{\text{sd}(\tilde{X})}$, (4)
which are the transformed (i.e., identified) set of estimates of the latent trait. In other words, the estimates have been scaled to have mean zero and standard deviation one within the sample. Consequently, in the latent case the ultimate estimates $\hat{X}$ have the same variance as the true (latent) regressor $X$, whereas in the standard case above, $\textrm{V}(\tilde{X})=\textrm{V}(X)+\textrm{V}(U)>\textrm{V}(X)$. This rescaling changes the attenuation factor in the latent case.
In the latent predictor case, we have $\sigma^{2}_{X}=1$ by identification and since our predictor when estimating our regression will be $\hat{X}$ which is equal to $\tilde{X}$ divided by $\sqrt{1+\sigma^{2}_{U}}$ (ignoring the mean adjustment which only affects the regression intercept), it follows that $\hat{\beta}_{\hat{X}}=\hat{\beta}_{\tilde{X}}\sqrt{1+\sigma^{2}_{U}}$ and thus
$\operatornamewithlimits{plim}\hat{\beta}_{\hat{X}}=\frac{1}{\sqrt{1+\sigma_{U}^{2}}}\beta_{X}.$ (5)
In other words, the slope coefficient estimate in the latent predictor case is still attenuated, but not as severely as in the standard case since $0<\frac{1}{1+\sigma_{U}^{2}}<\frac{1}{\sqrt{1+\sigma_{U}^{2}}}<1$ for $\sigma_{U}^{2}>0$.
In the next section, we review commonly used approaches for handling measurement error with latent predictors and show how several can worsen bias once identification rescaling is taken into account. We then introduce our split‑indicator correlation correction.
@@ 2.1 No Adjustment
The most common approach to dealing with measurement error in latent predictors has been to simply ignore the problem and use point estimates of the latent trait as the predictor without any further adjustment. In cases where measurement error is small relative to the variance of the predictor, using this standard approach may be reasonable. But with significant measurement error, slope estimates will be biased toward zero, sometimes dramatically so, when using standard regression estimates.
2.2 Full Joint Estimation: Benefits and Practical Costs
When attenuation from treating estimated latent scores as error-free is a concern, the most principled remedy is to estimate the measurement and outcome models jointly—treating $X$ as latent throughout—for example, with Bayesian factor–outcome models [undefy, undefe] or SEM/LISREL. This joint approach is the most principled way to run a regression with a latent predictor. But it is demanding in practice. It requires a full model for the indicators and the outcome, identification choices for scale and sign, and estimation of the joint posterior or likelihood (often via MCMC or SEM). This can be demanding and often infeasible for “black‑box” scores (e.g., PCA, additive indices, or ML/text embeddings) that lack an explicit likelihood (or one that can be tractably deployed via MCMC). As a consequence, researchers often employ other methods.
Because joint estimation is the gold-standard benchmark whenever it is feasible, we treat it as the reference point for evaluating the simpler, two-stage alternatives that we investigate. As we show, our proposed method can provide similar performance (e.g., RMSE) to full joint estimation, while other existing methods perform much worse.
2.3 Instrumental Variables Produce Over-Correction
When full joint estimation is infeasible, a standard errors-in-variables workaround is instrumental variables (IV), which aims to recover $\beta_{X}$ by isolating variation due to the signal rather than measurement noise (cf. [undefq]). In the classical setting with $\tilde{X}=X+U$, IV/2SLS corrects attenuation by using an instrument $Z$ that satisfies: (i) $\textrm{Cov}(X,Z)\neq 0$ (relevance); (ii) $\textrm{Cov}(Z,U)=0$; and (iii) $\textrm{Cov}(Z,\varepsilon)=0$ (exclusion). Under these conditions, 2SLS effectively leverages the part of the observed regressor that co-moves with ZZ—the component attributable to $X$, not UU—to identify the target slope. [^3] Although we focus on measurement error here, IV can also address simultaneity and omitted-variable bias under suitable assumptions (see [undefae]).
This logic is appealing in latent-variable applications because external instruments are often unavailable, but one can sometimes construct internal instruments by producing multiple estimates of the same latent trait from disjoint information. In the classical errors-in-variables case, this “two-measures” strategy can plausibly satisfy the IV conditions: the alternate measure is correlated with $X$ and, if built from independent information, can be approximately uncorrelated with the measurement error in the primary measure (and unrelated to $\varepsilon$).
[…]
This over-correction illustrates a broader theme: classical measurement-error fixes do not automatically carry over to latent predictors once identification rescaling is accounted for. In the next subsection, we show that the method of composition can also worsen bias, but in the opposite direction by exacerbating attenuation. We then introduce our split-indicator correlation correction, which can be viewed as an IV-style strategy that explicitly adjusts for identification-induced rescaling.